The thermal conductivity of copper is four times that of brass. Two rods of copper and brass having the same length and cross-section are joined in series.the free end of the copper rod is at 0 degree Celsius and that of the brass rod is at 100degree Celsius the temperature of the junction of the rod in equilibrium is

Grade:12th pass

2 Answers

Monu
4 years ago
Ans:Let the thermal conductivity of Cu be 4K and of Brass be K. 
Length of Copper and Brass rod ne L and their cross section area be A. 
Let the temperature at junction be T. 
By using formula:
KA(100-T) /L = 4KA(T-0) /L
100-T= 4T
5T=100
T=20°C.
 
 
Monu
4 years ago
Thermal conductivity of brass K of copper be 4K. Lengths be L and cross section area be A. 
Let the temperature at junction be X.  
Then, KA(100-X) /L = 4KA(X-0) /L
100-X=4X
5X=100
X= 20°C